# Taming Data Challenges in ML-based Security Tasks: Lessons from Integrating Generative AI

### v1

@@ VII-A Cases where GenAI shows potential
Traditional schemes (also sample-conditioned), SMOTE and MC-CCR perform poorly for this task. For 7 out of the 8 months, these schemes produce either negative gains or unreliable performance estimates. This highlights the limitations of traditional methods to mitigate concept drift. Failure scenarios. Two schemes GReaT and REaLTabFormer failed to converge on BODMAS. GReaT could not be trained on this task due to its feature dimensionality exceeding 1,000. After feature processing, GReaT needs to be trained on over 21K tokens per sample, which far exceeds the token limit 1024 of the backend LM. REaLTabFormer’s training failed as encoded features do not fit into the memory (NVIDIA A100 GPU).
@@
MS-Malware.    MS-Malware, a multi-class classification task (9 classes) suffers from limited or biased training data challenges (a), (b) and (d) (Section II), with the notable challenge being the high dimensionality of features (1,804 features). Results are in Table II.
@@
Class-conditioned GenAI schemes, including our new Nimai-C scheme, show significant performance improvement. Amongst the class-conditioned GenAI schemes, large standard deviation values make it difficult to directly compare means. We thus perform the Kruskal-Wallis test, and find no difference in rankings of these schemes due to a p-value of 0.286 $>$ 0.05. Interestingly, our sample-conditioned scheme, Nimai-S, despite achieving a performance improvement, does not perform as well as the class-conditioned schemes. Similar to the BODMAS task, we suspect that this task has test samples that have drifted from the training data distribution and requires data points in more novel/diverse regions of the data manifold to achieve better performance. Recall that class-conditioning is well suited for this goal. This is further confirmed by the fact that the traditional sample-conditioned schemes, SMOTE and MC-CCR perform poorly in this task.
@@
Why does augmentation improve performance?.    We perform a deep dive into IoT and BGP tasks to understand why synthetic data improved performance. Our hypothesis is that synthetic data helped mitigate certain biases in the real dataset, thereby enhancing generalization performance. We randomly selected 50 test instances per task that Nimai-S’s synthetic samples helped classify correctly after being misclassified before augmentation. Using Local Interpretable Model-agnostic Explanations (LIME) [80], we extract the top 10 most influential features for each corrected prediction and observe substantial distributional changes. We measure the distribution changes using entropy i.e., the randomness of the probability distribution; and skewness [30], a measure of the degree of asymmetry of a probability distribution. Skewness ranges from $\infty$- ∞ to +$\infty$+ ∞, indicating the dominant tail of a distribution: negative for lower values, positive for higher, and zero for symmetry. Classifiers generalize better on symmetric distributions.
For the IoT task, these features show a 36.8% increase in entropy and a 93.87% absolute reduction in skewness towards 0; for the BGP task, entropy increases by 175.33% and absolute reduction of skewness towards 0 by 72.8%. For both tasks, there is a consistent reduction in the absolute value of skewness for most features after augmentation, depicted in Figure 5 (Appendix). These shifts suggest that synthetic augmentation reduces bias and promotes more symmetric feature distributions, leading to better classifier generalization.
VII-B Cases where most GenAI schemes show insufficiency
@@
We study three tasks, Tor, Cookie, and nPrintML, where we discover significant challenges with improving performance using GenAI schemes.
Tor.   Tor is a binary classification task that suffers from limited or biased training data challenges (a)-(c) (Section II). The results are in Table IV.
@@
Finding 4.   Tasks with significant class overlap are challenging for GenAI-based data augmentation. No scheme except REaLTabFormer produces a positive gain (considering only reliable measurements). To understand this result, we visualize the t-SNE [102] feature representations of Tor samples and contrast them with the BGP samples in Figure 4. Recall that the GenAI schemes showed significant performance gains for the BGP task. We see that Tor classes appear largely inseparable even in their non-linearly projected feature spaces, i.e., with a large volume of class overlap, whereas BGP classes have distinct separability. This indicates that GenAI data augmentation faces challenges in datasets with class overlap, often caused by suboptimal features or noisy labels.
REaLTabFormer is the only method to demonstrate improvement on Tor task. This is likely due to its use of a statistical measure for early stopping, which helps avoid both overfitting and underfitting. Thus, the model can be properly instantiated even in the presence of overlapping class samples. Recall that REaLTabFormer failed in all previous settings, but here the lower feature dimensionality (175) and sufficient training samples also helped the model converge. The success of REaLTabFormer highlights the promise of correctly instantiated LLM-based GenAI schemes for tasks with class overlap.
@@
Failure cases. Only GReaT failed to converge on the Tor task. GReaT requires a large compute time to train on the Tor task—an estimated 22 hours per epoch (Appendix Table VII).
Figure 4: t-SNE plot for BGP and Tor task data.
@@
Cookie and nPrintML.    Cookie and nPrintML are multi-class classification tasks with 4 and 13 classes, respectively. The Cookie task has limited or biased data challenge (a) & (d), whereas nPrintML task has limited or biased data challenges (a), (b) and (d) (Section II). The results are shown in Table IV.
@@
1) Overlapping class distributions. Similar to Tor, we suspect the issue of overlapping class distributions. In the Cookie task, 3,707 samples in the training set have identical feature vectors assigned to more than one class label. The authors of this work acknowledge the issue of noisy labels because ground truth labels were obtained from multiple sources that may not agree uniformly on the labels. This is a clear problem. On the other hand, for nPrintML, while we are unable to find any evidence for noisy labels, the authors hint at poor distinguishability of samples among certain classes—suggesting potential class overlap. Figure 6(b) in their work [48], shows poor fingerprinting performance among samples belonging different variants of the same major OS, namely Windows, and Linux. The authors attribute this problem to the different variants (of the same OS) sharing similar network stack, leading to more similarity in the extracted feature vectors.
@@
  Nimai-S  -3.18 (0.00) $\downarrow$ -0.03 (0.66) - -0.16 (0.08) $\downarrow$
Nimai-C  -3.87 (0.00) $\downarrow$ -0.23 (0.57) - -0.1 (0.07) $\downarrow$
@@
TabSyn  -19.54 (7.96) $\downarrow$ $\times$ $\times$ 0.01 (0.08) -
@@
CTAB-GAN+  5.89 (12.52) - 0.3 (0.50) - -0.11 (0.07) $\downarrow$
TabDDPM  -21.75 (14.03) $\downarrow$ -0.22 (0.62) - -0.19 (0.08) $\downarrow$
@@
TABLE IV: Mean $\Delta G$ with standard deviation for Tor, nPrintML and Cookie tasks, where GenAI schemes do not show performance improvement. $\uparrow$ and $\downarrow$ indicate positive and negative gains, respectively. ‘-’ indicates unreliable gains (CV $\geq$ 1), and ‘$\times$’ indicates model failures.
@@
2) Sparse feature vectors. The training datasets of nPrintML and Cookie show high feature sparsity (percentage of feature values that are zero)—78.73% and 98.66%, respectively. To assess the impact of sparsity on performance, we run a new experiment, minimizing potential class overlap by reducing the 13-class nPrintML task to 5 carefully selected classes.. [^3] We only retain 5 OS classes - i.e. Mac, Kali-linux, Ubuntu-desktop (combining all variants), Ubuntu-server and Windows (combining all variants). SMOTE applied to this setting produced a mean $\Delta G$ of -0.31%, showing no improvement in performance, suggesting sparsity plays a role in the non-improvement of performance.
To further examine whether GenAI schemes struggle with performance improvement in noisy/overlapping data, we artificially induce class overlap in BGP. Recall that GenAI schemes (Nimai-S and CTAB-GAN+) performed well on BGP. We flip 20% of class 0 samples to class 1 to induce overlap. This resulted in these GenAI methods yielding negative gains of up to -37%. This further strengthens our Finding 5. The experimental setup and results are detailed in Appendix Section -H.
@@
Failure scenarios. Four GenAI schemes, GReaT, REaLTabFormer, TVAE and TabSyn failed in one or both tasks. We were unable to train GReaT for both tasks. Both tasks have feature dimensionality over 1,000, resulting in over 11K and 28K tokens per sample, which is inordinately greater than the token limit of 1024 of the backend LLMs. REaLTabFormer failed to train on both tasks. For nPrintML, training fails due to large encoded features which does not fit into the memory (NVIDIA A100 GPU). For Cookie, training does not converge as our early stopping criteria is not met even after 40 hours. TVAE failed for the Cookie task, because we failed to generate new samples, even after 100 hours of rejection sampling (see Appendix -C). TabSyn failed for the nPrintML task. TabSyn’s VAE training on the pre-processed data does not fit into the memory (NVIDIA A100 GPU) for nPrintML due to the large feature size and number of classes.
@@
TABLE V: Mean macro F-scores and standard deviation for Nimai-C and Nimai-hybrid across 10 trials and macro F-scores for w/o GenAI and INSOMNIA on BODMAS months 5 & 6. (-) indicates no standard deviation.
VIII Evaluation: Post-Deployment
GenAI schemes are still limited by existing training data in improving classifier generalization. We again focus on a concept drift setting, where GenAI schemes may struggle to generate points that fit the drifted test. To mitigate drift after deployment, the traditional practice involves a costly and time-consuming labeling process for the drifted test set, followed by retraining the classifier on the new data. Security threats are known to change abruptly [6], requiring a faster reaction (e.g., for NIDS and malware detection). In this section, we investigate whether GenAI can be used to perform faster recovery from concept drift, without requiring a complete data labeling effort on the test set. To achieve this, we propose a novel scheme called Nimai-hybrid that combines both the Nimai-C and Nimai-S approaches.
Similar to the setup in Section VII-A, we train our Nimai scheme on month 0 data. We focus to mitigate concept drift on months 5 and 6, which show the highest performance drop (Figure 2). We assume the defender has detected the concept drift in month $k$ ($k\in[1,5]$). Existing methods such as CADE [111] and Transcend [51, 10] can be used to detect concept drift. To rapidly recover from concept drift, we use Nimai-S’s sample conditioning to generate synthetic points from a small labeled subset of the drifted test set from month $k$. This is possible as Nimai can perform sample-conditioned generation on new samples without the need for retraining, unlike other GenAI techniques that perform class-conditioned generation. The samples for labeling are chosen using uncertanity sampling [6], an active learning technique [4]. It chooses samples close to the decision boundary based on classifier confidence scores. We show that even a small subset is sufficient to boost performance via data augmentation, with subset sizes ranging from 9 to 64 for the BODMAS task (see Table X). This significantly reduces the cost of labeling effort, speeding adaptation to drift, i.e., fast recovery.
For a testing month $N$ ($N$ can be months 5 and 6), we evaluate whether uncertainty samples from any of the previous $(N-1)$ months can improve performance. We use a confidence interval of 0.3-0.7 from a random 1% subset of the test set to determine uncertainty samples. We augment Nimai-S with 5 times the synthetic samples conditioned on uncertain samples. The remaining synthetic samples are generated using Nimai-C (i.e., class-conditioned) to balance the class sizes. We compare Nimai-hybrid with the three baseline schemes: (1) w/o GenAI: This reports classifier performance without any data augmentation, i.e., only using the real data. (2) Nimai-C: This mirrors the setting in Section VII-A, where class conditioning uses only month 0 data for generating synthetic samples. (3) INSOMNIA: The security classifier is retrained on a combination of real and uncertainty samples (the same samples used by Nimai-hybrid). This strategy is equivalent to the existing method called INSOMNIA [6] which aims to adapt classifiers to concept drift using uncertainty samples. We use the macro F-score, reporting the mean and standard deviation across 10 trials for Nimai-C and Nimai-hybrid.
@@ VIII Evaluation: Post-Deployment
Finding 6.   A combination of class and sample-conditioned GenAI schemes can help to quickly recover from concept drift with low-cost labeling effort. For months 5 and 6, we present the best performance obtained by any of the previous $(N-1)$ months across 10 trials of data augmentation. Results for months 5 and 6 are shown in Table V. Complete results are in Table X (Appendix). Due to high standard deviations, we applied Kruskal-Wallis test to compute ranks in both months. In month 5, Nimai-hybrid and Nimai-C show no difference in means, making them equivalent schemes. In month 6, which shows the maximum drop in performance in all test months of BODMAS, we find Nimai-hybrid outperforms Nimai-C. This shows the potential for Nimai to rapidly recover using drifting samples post-deployment. We conducted our experiment for the remaining 10 months (which did not degrade as much as months 5 and 6) but observed no performance improvement from using Nimai-hybrid, indicating room for advancing our recovery strategy. However, we provided significant gains for the worst affected months.
IX Conclusion & Future Work

### v2

@@ 7.1. Cases where GenAI shows potential
Failure scenarios. Two schemes GReaT and REaLTabFormer failed to converge on BODMAS. GReaT could not be trained on this task due to its feature dimensionality exceeding 1,000. After feature processing, GReaT needs to be trained on over 21K tokens per sample, which far exceeds the token limit 1024 of the backend LM. REaLTabFormer’s training failed as encoded features do not fit into the memory (NVIDIA A100 GPU).
@@
MS-Malware.    MS-Malware, a multi-class classification task (9 classes) suffers from limited or biased training data challenges (a), (b) and (d) (Section 2), with the notable challenge being the high dimensionality of features (1,804 features). Results are in Table 2.
@@
Class-conditioned GenAI schemes, including our new Nimai-C scheme, show significant performance improvement. Amongst the class-conditioned GenAI schemes, large standard deviation values make it difficult to directly compare means. We thus perform the Kruskal-Wallis test, and find no difference in rankings of these schemes due to a p-value of 0.286 >> 0.05. Interestingly, our sample-conditioned scheme, Nimai-S, despite achieving a performance improvement, does not perform as well as the class-conditioned schemes. Similar to the BODMAS task, we suspect that this task has test samples that have drifted from the training data distribution and requires data points in more novel/diverse regions of the data manifold to achieve better performance. Recall that class-conditioning is well suited for this goal, as confirmed by the poor performance of traditional sample-conditioned schemes such as SMOTE and MC-CCR.
@@
Why does augmentation improve performance?.    We perform a deep dive into IoT and BGP tasks to understand why synthetic data improved performance. Our hypothesis is that synthetic data helped mitigate certain biases in the real dataset, thereby enhancing generalization performance. We randomly selected 50 test instances per task that Nimai-S’s synthetic samples helped classify correctly after being misclassified before augmentation. Using Local Interpretable Model-agnostic Explanations (LIME) (ribeiro2016should), we extract the top 10 most influential features for each corrected prediction and observe substantial distributional changes. We measure the distribution changes using entropy i.e., the randomness of the probability distribution; and skewness (doane2011measuring), a measure of the degree of asymmetry of a probability distribution. Skewness ranges from -$\infty$ to $\infty$, indicating the dominant tail of a distribution: negative for lower values, positive for higher, and zero for symmetry. Classifiers generalize better on symmetric distributions. For the IoT task, these features show a 36.8% increase in entropy and a 93.87% absolute reduction in skewness towards 0; for the BGP task, entropy increases by 175.33% and absolute reduction of skewness towards 0 by 72.8%. For both tasks, there is a consistent reduction in the absolute value of skewness for most features after augmentation, depicted in Figure 5 (Appendix). These shifts suggest that synthetic augmentation reduces bias and promotes more symmetric feature distributions, leading to better classifier generalization.
7.2. Cases where most GenAI schemes show insufficiency
@@
We study 3 tasks, Tor, Cookie, and nPrintML, where we discover significant challenges to improve performance using GenAI schemes.
@@
Tor.   Tor is a binary classification task that suffers from limited or biased training data challenges (a)-(c) (Section 2). Table 4 reports the results. Finding 4.   Tasks with significant class overlap are challenging for GenAI-based data augmentation. No scheme except REaLTabFormer produces a positive gain (considering only reliable measurements). To understand this result, we visualize the t-SNE (van2008visualizing) feature representations of Tor samples and contrast them with the BGP samples in Figure 4. Recall that the GenAI schemes showed significant performance gains for the BGP task. We see that Tor classes appear largely inseparable even in their non-linearly projected feature spaces, i.e., with a large volume of class overlap, whereas BGP classes have distinct separability. This indicates that GenAI data augmentation faces challenges in datasets with class overlap, often caused by suboptimal features or noisy labels. REaLTabFormer is the only method to improve on the Tor task, likely due to its statistical early-stopping criterion that prevents over- and underfitting. Unlike prior settings where it failed, the lower feature dimensionality (175) and ample training samples here enabled convergence. The success of REaLTabFormer highlights the promise of correctly instantiated LLM-based GenAI schemes for tasks with class overlap.
Figure 4. t-SNE plot for BGP and Tor task data. :
@@
Failure cases. Only GReaT failed to converge on the Tor task. GReaT requires a large compute time to train on the Tor task—an estimated 22 hours per epoch (Appendix 8 (Appendix)).
@@
Cookie and nPrintML.    Cookie and nPrintML are multi-class tasks with 4 and 13 classes, respectively. Cookie faces challenges (a) and (d), while nPrintML faces (a), (b), and (d) (Section 2). Results are in Table 4.
@@
1) Overlapping class distributions. Similar to Tor, we suspect the issue of overlapping class distributions. In the Cookie task, 3,707 samples in the training set have identical feature vectors assigned to more than one class label. The authors of this work acknowledge the issue of noisy labels because ground truth labels were obtained from multiple sources that may not agree uniformly on the labels. This is a clear problem. On the other hand, for nPrintML, while we are unable to find any evidence for noisy labels, the authors hint at poor distinguishability of samples among certain classes—suggesting potential class overlap. Figure 6(b) in the original work (holland2021new), shows poor fingerprinting performance among samples belonging different variants of the same major OS, namely Windows, and Linux. The authors attribute this problem to the different variants (of the same OS) sharing similar network stack, leading to more similarity in the extracted feature vectors.
@@
  Nimai-S  -3.18 (0.00) $\downarrow$ 0.03 (0.66) - -0.06 (0.17) $\downarrow$
Nimai-C  -3.87 (0.00) $\downarrow$ -0.23 (0.57) - -0.05 (0.11) $\downarrow$
@@
TabSyn  -19.54 (7.96) $\downarrow$ $\times$ $\times$ 0.11 (0.18) -
@@
CTAB-GAN+  5.89 (12.52) - 0.3 (0.50) - 0.02 (0.14) -
TabDDPM  -21.75 (14.03) $\downarrow$ -0.22 (0.62) - -0.01 (0.11) $\downarrow$
@@
Table 4. Mean $\Delta G$ with standard deviation for Tor, nPrintML and Cookie tasks, where GenAI schemes do not show performance improvement. $\uparrow$ and $\downarrow$ indicate positive and negative gains, respectively. ‘-’ indicates unreliable gains (CV $\geq$ 1), and $\times$’ indicates model failures.
@@
2) Sparse feature vectors. The training datasets of nPrintML and Cookie show high feature sparsity (percentage of feature values that are zero)—78.73% and 98.66%, respectively. To assess the impact of sparsity on performance, we run a new experiment, minimizing potential class overlap by reducing the 13-class nPrintML task to 5 carefully selected classes. [^4] We only retain 5 OS classes - i.e. Mac, Kali-linux, Ubuntu-desktop (combining all variants), Ubuntu-server and Windows (combining all variants). SMOTE applied to this setting produced a mean $\Delta G$ of -0.31%, showing no improvement, suggesting sparsity plays a role in the non-improvement of performance. To further examine whether GenAI schemes struggle with performance improvement in noisy/overlapping data, we artificially induce class overlap in BGP. Recall that GenAI schemes (Nimai-S and CTAB-GAN+) performed well on BGP. We flip 20% of class 0 samples to class 1 to induce overlap. This resulted in these GenAI methods yielding negative gains of up to -37%. This further strengthens our Finding 5. More details are in Appendix H.
@@
Failure scenarios. Four GenAI schemes, GReaT, REaLTabFormer, TVAE and TabSyn failed in one or both tasks. We were unable to train GReaT for both tasks. Both tasks have feature dimensionality over 1,000, resulting in over 11K and 28K tokens per sample, which is inordinately greater than the token limit of 1024 of the backend LLMs. REaLTabFormer failed to train on both tasks. For nPrintML, training fails due to large encoded features which does not fit into the memory (NVIDIA A100 GPU). For Cookie, training does not converge as our early stopping criteria is not met even after 40 hours. TVAE failed for the Cookie task, because we failed to generate new samples, even after 100 hours of rejection sampling (see Appendix C). TabSyn failed for the nPrintML task. TabSyn ’s VAE fails to fit in GPU memory (NVIDIA A100) for nPrintML due to its large feature size and many classes.
8. Evaluation: Post-Deployment
GenAI schemes are still limited by existing training data in improving classifier generalization. We again focus on a concept drift setting, where GenAI schemes may struggle to generate points that fit the drifted test. To mitigate drift after deployment, the traditional practice involves a costly and time-consuming labeling process for the drifted test set, followed by retraining the classifier on the new data. Security threats are known to change abruptly (andresini2021insomnia), requiring a faster reaction (e.g., for NIDS and malware detection). In this section, we investigate whether GenAI can be used to perform faster recovery from concept drift, without requiring a complete data labeling effort on the test set. To achieve this, we propose a novel scheme called Nimai-hybrid that combines both the Nimai-C and Nimai-S approaches.
Similar to the setup in Section 7.1, we train our Nimai scheme on month 0 data. We focus to mitigate concept drift on months 5 and 6, which show the highest performance drop (Figure 2). We assume the defender has detected the concept drift in month $k$ ($k\in[1,5]$). Existing methods such as CADE (yang2021cade) and Transcend (jordaney2017transcend; barbero2022transcending) can be used to detect concept drift. To rapidly recover from concept drift, we use Nimai-S’s sample conditioning to generate synthetic points from a small labeled subset of the drifted test set from month $k$. This is possible as Nimai can perform sample-conditioned generation on new samples without the need for retraining, unlike other GenAI techniques that perform class-conditioned generation. The samples for labeling are chosen using uncertanity sampling (andresini2021insomnia), an active learning technique (aggarwal2014active). It chooses samples close to the decision boundary based on classifier confidence scores. We show that even a small subset is sufficient to boost performance via data augmentation, with subset sizes ranging from 9 to 64 for the BODMAS task. This significantly reduces the cost of labeling effort, speeding adaptation to drift, i.e., fast recovery.
For a testing month NN (NN can be months 5 and 6), we evaluate whether uncertainty samples from any of the previous (N1)(N-1) months can improve performance. We use a confidence interval of 0.3-0.7 from a random 1% subset of the test set to determine uncertainty samples. We augment Nimai-S with 5 times the synthetic samples conditioned on uncertain samples. The remaining synthetic samples are generated using Nimai-C (i.e., class-conditioned) to balance the class sizes.
We compare Nimai-hybrid with the three baseline schemes: (1) w/o GenAI: This reports classifier performance without any data augmentation, i.e., only using the real data. (2) Nimai-C: This mirrors the setting in Section 7.1, where class conditioning uses only month 0 data for generating synthetic samples. (3) INSOMNIA: The security classifier is retrained on a combination of real and uncertainty samples (the same samples used by Nimai-hybrid). This strategy is equivalent to the existing method called INSOMNIA (andresini2021insomnia) which aims to adapt classifiers to concept drift using uncertainty samples. We use the macro F-score, reporting the mean and standard deviation across 10 trials for Nimai-C and Nimai-hybrid.
@@ 8. Evaluation: Post-Deployment
Table 5. Mean macro F-scores and standard deviation for Nimai-C and Nimai-hybrid across 10 trials and macro F-scores for w/o GenAI and INSOMNIA on BODMAS months 5 & 6. (-) indicates no standard deviation.
@@
Finding 6.   A combination of class and sample-conditioned GenAI schemes can help to quickly recover from concept drift with low-cost labeling effort. For months 5 and 6, we present the best performance obtained by any of the previous (N1)(N-1) months across 10 trials of data augmentation. Results for months 5 and 6 are shown in Table 5. Complete results are in Table 11 (Appendix). High deviations led us to use the Kruskal–Wallis test for ranking in both months. In month 5, Nimai-hybrid and Nimai-C show no difference in means, making them equivalent schemes. In month 6, which shows the maximum drop in performance in all test months of BODMAS, we find Nimai-hybrid outperforms Nimai-C. This shows the potential for Nimai to rapidly recover using drifting samples post-deployment. We conducted our experiment for the remaining 10 months (which did not degrade as much as months 5 and 6) but observed no performance improvement from using Nimai-hybrid, indicating room for advancing our recovery strategy. However, we provided significant gains for the worst affected months.
9. Conclusion & Future Work